1) Describe how you demonstrate that white light is compose of a number of colours.
2) give an example of magnetic induction and mention the feature which distinguish magnetic induction from other process of magnetization.
3) describe two example of resonance
4) define wavelength,frequency and wave velocity and also deduce a relationship between them.
1
Expert's answer
2019-08-27T09:35:31-0400

"1." A white light when passes through a triangular glass prism gets disperse into seven colour

Morever,

A Newton's disc can be created by painting a disc with the seven different colours":" violet, indigo, blue, green, yellow, orange and red. When the disc is rotated, it appears white. This explains that white light consists of seven colours.

2"." Example"-" When an iron needle is attached to a magnet magnetism induces in it . Therefore on acquiring its magnetism it acquires the magnetic poles.


Feature which distinguish magnetic induction from other process of magnetization are-

"\to" In magnetic induction there is not any stroke by one end of the magnet from one end to other end.

"\to"It is temporary magnetism. materials is magnetises as long as it is attached to pole of magnet.

"\to" There is not any electric current required as in the case of electrical magnetism.


"3."

(a) A familiar example is a playground swing, which acts as a pendulum. Pushing a person in a swing in time with the natural interval of the swing (its resonant frequency) makes the swing go higher and higher (maximum amplitude), while attempts to push the swing at a faster or slower tempo produce smaller arcs.


(b)Shattering of a crystal wineglass when exposed to a musical tone of the right pitch (its resonant frequency).

"4."Wavelength"-" One wavelength equals the distance between two successive wave crests or troughs.

It is denoted by"\lambda" .

Frequency"-"It is defined as the number of waves that passes a given point per second.It is denoted by "\nu".

Wave Velocity "-"  It is the velocity at which a wave phase  propagates at a certain frequency. It is denoted by "c" .The velocity of a wave is equal to the product of its wavelength and frequency (number of vibrations per second) and is independent of its intensity.

Relation between "\lambda,\nu" and "c" :

Velocity"=" "\frac{wavelength}{timeperiod}" "=wavelength\times{frequency}"

"So,
c=\nu\times\lambda"